OpenStudy (anonymous):

im on flvs and i need to get this assignment done fast so i can bring up my grade can someone help

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

well can you

OpenStudy (anonymous):

what is the question???

OpenStudy (anonymous):

Assessment 02.04 Module Two Quiz Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem. A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

is this 4 questions???

OpenStudy (anonymous):

yeah 1 sec

OpenStudy (anonymous):

this is it